N2 effuses at a rate of 4.5 m/s through a specific system. What is the rate of effusion (m/s) for CH4 in the same system?
Answer #1

The calculations for the rate of diffusion is given as follows

According to Grahams Law of Effusion, the rate of effusion of a gas is inversely proportional to the square root of its molar mass. So,

rate gas1/rate gas2 = sq.rt. MW2 / sq.rt. MW1

Since molar mass CH4 is less than N2, it will effuse faster

N2 = 28 and CH4 = 16

4.5 / x = sq.rt. (16 / 28)

4.5 / x = 4 / 5.29

4.5/x = 0.756

x = 5.95 m/s = 6 m/sec

Thus, the rate of effusion (m/s) for CH4 in the same system would be 6 m/s